# Unraveling the Enigma of Double Descent: An In-depth Analysis through the Lens of Learned Feature Space

### v1

@@ 4.1 MNIST
Figure 1: The phenomenon of double descent on two-layer fully connected neural networks trained on MNIST $(N=4000)$, under varying label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labelled data denoted as ${\rm Kp}$ when $p>0$. Optimized with SGD for 4000 epochs and decreasing learning rate. The test error curve of $p=[10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy curve ${\rm Kp}$ of in-context noisy data learning showed a correlation generalization performance beyond context.
@@
We commence our exploration by revisiting one of the earliest experiments that revealed the presence of the double descent phenomenon, as presented in the work of Belkin et al. (2019). This experiment involved the training of a basic two-layer fully connected neural network on a dataset comprising 4000 samples from the MNIST dataset ($N=4000$). When zero label noise ($p=0\%$) exists, as depicted in Figure.1(a), both the train and the test error curve decreases monotonically with increasing model width. The training error exhibits a faster decline and approaches zero at approximately $k=15$, indicating the occurrence of an interpolation threshold under these conditions. The test error still decreases after this interpolation threshold, demonstrated a benign over-parameterization.
@@
With label noise ($p=10\%/20\%$) introduced in the training dataset, a test error peak starts to form around the interpolation threshold as showed in Figure.1(b, c). Starting with small models, the test error curve exhibits a U-shaped pattern in the first half and the bias-variance trade-off is reached at the bottom of the U-shaped curve. The decrease of both train and test error before this bias-variance trade-off indicates the model undergoes an under-fitting stage. With the expansion of the model size, the model transitions into an overfitting regime, leading to a peak in the test error around the interpolation threshold, while the train error converges to 0. Finally, after surpassing this test error peak, the testing loss experiences a decline once more, showcasing how highly over-parameterized models can ultimately yield improved generalization performance. This intriguing behavior is referred to as double descent. As the test loss undergoes a pattern of first decreasing, then increasing, and finally decreasing again, the test accuracy exhibits the opposite trend, i.e., first increasing, then decreasing, and then increasing once more. As Nakkiran et al. reported, the test loss peak is positively correlated with the label noise ratio $p$. In other words, as the label noise ratio increases, the peak in the test loss also tends to increase accordingly.
@@
To comprehend the influence of noisy labeled data on the variation of the learned model during the double descent phenomenon, we examine the learned representations separately for the class of clean and noisy labeled data, presented as the prediction accuracy ${\rm Kp}$ of noisy labelled data by their 5-nearest neighbour of clean labelled data. In this experiment setup, as depicted in Figure.1(b, c), the ${\rm Kp}$ curve follows a opposite but similar trend to the test loss curve. When test loss first decreases as model expands, the accuracy ${\rm Kp}$ rises to about 60% and then decreases around the threshold where the peak of test error arises. Then, as the test error reduces and test accuracy rises, the ${\rm Kp}$ curve rises again to above 90% for highly over-parameterized models. Recall that the noisy data points will be predicted based on their randomly assigned labels during this stage as the model fully interpolates the training set, we interpret the ascending ${\rm Kp}$ curve in the over-parameterized region as a valid ‘isolation’ of the noisy labeled data in the learned feature space. When the neural network positions the learned representation of noisy data points closer to the class associated with their ‘incorrect’ labels, it results in a lower prediction accuracy, denoted as ${\rm Kp}$. Under this circumstances, similar test samples also tend to be predicted incorrectly due to the influence of these noisy labels. If the neural network successfully identifies the underlying data patterns and situates the learned representations of noisy data points closer to their correct class, the prediction accuracy, denoted as ${\rm Kp}$, will be high by nearby class members. Consequently, test samples resembling the noisy training data have a higher likelihood of being accurately identified due to their proximity to the surrounding clean data points, and becoming less susceptible to the influence of the ‘isolated’ noisy labeled data. This explanation accounts for the improved generalization performance for over-parameterized models.
Figure 2: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained SimpleFC of 1000 randomly selected data points from the training set of MNIST, at different layer widths of $k=[10,20,100]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models are able to segregate noise within the learned feature space.
The t-SNE visualization in Figure.2 offers evidence from a more intuitive standpoint: as layer width increases, the overlap between different classes diminishes, and noisy data points are more accurately assigned to their correct classes. The choice of layer width $k=[10,20,100]$ represents three different circumstances. When $k=10$, bias-variance trade-off is reached at the bottom of the U-shaped segment of the test error curve. The training dataset is not fully interpolated and the learned representations are less organized. When $k=20$, the model are about to interpolate the whole training set, which includes the noisy labelled data. From the figure, it’s evident that at this stage, representations of some data points with noisy labels appear within classes they don’t belong to, which are likely to introduce interference and lead to misleading effects. In the case of an over-parameterized model, such as when $k=100$, this phenomenon significantly diminishes, and the majority of noisy labeled data points are encoded alongside their true classmates.
@@ 4.2 CIFAR-10
Figure 3: The phenomenon of double descent on five-layer CNNs trained on CIFAR-10 ($N=50000$), under varying label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labelled data denoted as ${\rm Kp}$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve of $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy curve ${\rm Kp}$ of in-context noisy data learning showed a correlation generalization performance beyond context.
Figure 4: T-Distributed Stochastic Neighbor Embedding (T-SNE) Visualization of the learned representation vectors by trained CNNs of 1000 randomly selected data points from the training set of CIFAR-10, at different layer widths of $k=[6,12,36]$. Circular points denote the clean data points, while asterisk points represent the data points with noisy labels (color-coded according to their original class). With an increase in layer width, the overlap between distinct classes diminishes, leading to a more precise allocation of noisy data to their respective correct classes, which indicates that over-parameterized models are able to segregate noise within the learned feature space.
@@
After a baseline case of deep learning on MNIST, we extended the experiment result to image classification on CIFAR-10, using two famous neural network architectures: CNN and ResNet. We first visit the experiment when the complete CIFAR-10 dataset ($N=50000$) with zero label noise ($p=0\%$) in trained with five-layer CNNs: we observe that in Figure.3(a), a proposed double descent peak arises around the interpolation threshold. We suppose that this phenomenon may arise from the intrinsic noise and increased complexity associated with image recognition of CIFAR-10 in contrast to MNIST. However this observation differentiates from our last experiment, we can still see from Figure.3(b,c) that when additional label noise ($p=10\%/20\%$) is introduced, the test error peak still increases accordingly, proving the positive correlation between noise and double descent. While label noise increases the difficulties of models fitting the train set and shifts the interpolation threshold rightwards, the test error peak also shifts correspondingly.
@@
When we review the noisy labelled data prediction test in the learned feature space, we can still observe a correlation between the trend of generalization performance and prediction accuracy ${\rm Kp}$. When test accuracy first increases then decreases as model expands, the accuracy ${\rm Kp}$ rises to 60% and then decreases before the threshold when the peak of test error arises. Then, as the test error reduces and test accuracy rises, the ${\rm Kp}$ curve rises again to 60% for over-parameterized models. The observed correlation between the trend of test loss and ${\rm Kp}$ also remains consistent. The same conclusion can be derived from the experiment context Figure.5 of ResNet18s trained on the sample task. In conclusion, by constructing ${\rm Kp}$ to represent the distribution of noisy labeled training data in the learned feature space, we propose that this distribution have a robust correlation to the model’s generalization performance across different levels of model width. We posit that this correlation can be attributed to the fact that the model’s mapping of noisy training data within the learned feature space is able to reflect the model’s prediction ability.
When we examine the t-SNE visualization of learned representations in this scenario, despite the increased complexity of 2D distributions due to the higher dimensionality of CIFAR-10 images, we can still observe distinct class distributions and improved identification of noisy data points within these distributions as the model width expands. The t-SNE visualization Figure.9 provided in A conveys the same statement.
@@
Figure 5: The phenomenon of double descent on ResNet18s trained on CIFAR-10 $(N=50000)$, under varying label noise of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labelled data denoted as ${\rm Kp}$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve under $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy curve ${\rm Kp}$ of in-context noisy data learning showed a correlation to the generalization performance beyond context.
@@ 5 Conclusion
In this paper, we have studied the feature space of learned representations for both clean and noisy training data points. By analyzing the behavior of the model when achieving a perfect fit on the training set, we observe that: as the model’s complexity and capacity increase, there is a clear and distinctive isolation of noisy data points within their original ‘class-mates’ in the learned feature space. This intriguing observation indicates that the neural network becomes increasingly capable of discerning and distinguishing noisy data from clean data with increasing model width, and plays a significant role in understanding the mechanism of deep learning generalization. As the model’s width expands, it gains a heightened sensitivity to the nuances and irregularities presented in the noisy data, allowing it to make finer differentiation between various classes. We posit that the mechanism behind this phenomenon can be attributed to two key factors. First, the gradient descent optimization algorithm endeavors to strike a perfect balance among training data points to minimize the loss function, which can lead to the isolation of noisy data points. Second, the characteristics of high-dimensional space may facilitate the beneficial separation of noisy data points from clean data in the learned feature space.
From the perspective of this paper, the observed double decent phenomenon is associated with imperfect models, first behaving as predicted by the classical learning theory and increasing their error on the test set until they reach full interpolation on the train set. Then, they make use of the extra parameters above the interpolation point to better separate noise from information. We intentionally referred to ‘imperfect models’ because we claim that double descent should not be observable in perfect learner models where perfect regularization should eliminate the effect. From our perspective, the double descent phenomenon is strictly related to imperfect models learning from noisy data.
Let’s expand on this by restricting our considerations to supervised learning for classification problems: For an ideal trading set with full information (no noise and no missing cases), a learner will learn the exact model at the interpolation point where the train set is exactly classified and the error becomes zero. In this case, at interpolation, also the test set will be exactly classified because the learner has learned the exact map. Before that interpolation, the error decreases with every new example. Clearly, this system does not show any double descent. At the interpolation point, the number of model parameters is approximately equal to the number of examples in the complete train set (i.e. the model is the matching table).
In the presence of noise, the learner cannot learn the full map by interpolation. Indeed, the bias-variance trade-off predicts a U-shaped curve in performances with first the learner increasing out-of-sample performances from digesting examples and then losing performances by trying to fit the in-sample noisy data. However, a learner who makes good use of regularization can do better and learn how to account for uncertainty. It is likely that the number of model parameters needed to learn both the underlying map and to handle the noise is larger than the number of parameters needed to learn only the map in the ideal case without noise (the interpolation point). If the model is a perfect learner with perfect regularization, we do not expect the manifestation of any double descent, but rather a gradual improvement in out-of-sample performances with the number of parameters until a saturation point is reached where performances cannot be improved without further examples. We expect that this saturation point should occur at a number of parameters larger than the interpolation point. Double descent is instead observed for imperfect learners that at first try to interpolate the noise and then start an intrinsic regularization via over-parameterization. Our particular model of the emergence of the double descent phenomenon is illustrated In Fig. 6.
@@
Figure 6: Exemplification of the emergence of the double descent effect in the perspective of the present paper. Starting from the bottom of the left panel: 1. the perfect learner, in the absence of noise, will learn the perfect model containing a number of parameters approximately equal to the number of significant examples; 2. in the presence of noise, the perfect learner needs to learn also to distinguish information from noise and it will learn a perfect model with a number of parameters which is higher the number of noisy examples; 3. the imperfect learner with optimal regularization will learn a well-performing model which distinguishes information from noise using a number of parameters which is higher the number of examples but should not show double descent peak. On the right panel: 4. the imperfect learner with sub-optimal regularization will learn a well-performing model first interpolating noise and signal and then learning to distinguish information from noise using a number of parameters which is higher the number of examples showing the double descent peak; 5. the larger is the noise, the higher the double descent peak is, and the larger is the number of parameters needed to distinguish information from noise.
@@
Reproducibility Statement   To facilitate reproducibility, we provide a detailed description of the neural architectures in Section 3 and full training details including exact settings of hyper-parameters in Appendix A. Our code is available at https://github.com/Yufei-Gu-451/double_descent_inference.
@@ Acknowledgments
The authors acknowledge the discussion with Siddharth Chaudhary, Joe Down, Nathan D’Souza, and Huijie Yan, all UCL students who took part in the academic year 2022-23 to project COMP0031. With them, we had invaluable discussions and great inspiration. This paper originated from that project and would not have been conceived without it. Please see Chaudhary et al. (2023) for the reference to that work. The author, T.A., acknowledges partial financial support from ESRC (ES/K002309/1), EPSRC (EP/P031730/1), and EC (H2020-ICT-2018-2 825215). X.Z. knowledges partial support from National Natural Science Foundation of China (No. 62076068).
@@ A.1 Scaling of Model Parameterization
The scaling of model parameterization to the universal layer width unit $K$ is showed in Figure.7.
@@
Figure 7: Scaling of number of parameters as model size with layer width unit k of the three neural architectures used in our experiments. We apply a logarithmic scale to the parameter counts $P$ of all neural architectures.

### v2

@@ 4.1 MNIST
Figure 1: The phenomenon of double descent on two-layer fully connected neural networks trained on MNIST $(N=4000)$, under varying explicit label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labeled data denoted as $P$ when $p>0$. Optimized with SGD for 4000 epochs and decreasing learning rate. The test error curve of $p=[10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation generalization performance beyond context.
@@
We commence our exploration by revisiting one of the earliest experiments that revealed the presence of the double descent phenomenon, as presented in the work of Belkin et al. (2019). This experiment involved the training of a basic two-layer fully connected neural network on a dataset comprising 4000 samples from the MNIST dataset ($N=4000$). When zero label noise ($p=0\%$) is introduced, as depicted in Figure.1(a), both the train and the test error curve decrease monotonically with increasing model width. The training error exhibits a faster decline and approaches zero at approximately $k=15$, indicating the occurrence of an interpolation threshold under these conditions. The test error still decreases after this interpolation threshold, demonstrating a benign over-parameterization. This noticeable alteration in the shape of the test error curve at the interpolation threshold may be attributed to the intrinsic label errors of MNIST (Northcutt et al., 2021a).
@@
With label noise ($p=10\%/20\%$) introduced in the training dataset, a more distinguishable test error peak starts to form around the interpolation threshold as shown in Figure.1(b, c). Starting with small models, the test error curve exhibits a U-shaped pattern in the first half and the bias-variance trade-off is reached at the bottom of the U-shaped curve. The decrease of both train and test error before this bias-variance trade-off indicates the model undergoes an under-fitting stage. With the expansion of the model size, the model transitions into an overfitting regime, leading to a peak in the test error around the interpolation threshold, while the train error converges to 0. Finally, after surpassing this test error peak, the testing loss experiences a decline once more, showcasing how highly over-parameterized models can ultimately yield improved generalization performance. This intriguing behavior is referred to as double descent. As the test loss undergoes a pattern of first decreasing, then increasing, and finally decreasing again, the test accuracy exhibits the opposite trend, i.e., first increasing, then decreasing, and then increasing once more. As Nakkiran et al. (2021) reported, the test loss peak is positively correlated with the label noise ratio $p$. In other words, as the label noise ratio increases, the peak in the test loss also tends to increase accordingly.
To comprehend the influence of noisy labeled data on the variation of the learned model during the double descent phenomenon, we examine the percentage $P$ of noisy data, and the majority of its nearest neighbours in feature space are in the same class. In this experiment setup, as depicted in Figure.1(b, c), the trajectory of the $P$ curve is in accordance with the pattern observed in the test accuracy curve, indicating a consistent alignment between the two. When test loss first decreases as the model expands, test accuracy and $P$ both rise. With the introduction of label noise, test accuracy decreases while $P$ fluctuates around 70%. When passing the interpolation threshold, both the test accuracy and $P$ significantly rise to 85% and nearly 100%, substantially affirming the validity of our hypothesis. In the meantime, the trend of test loss is opposite to both test accuracy and $P$, while changing correspondingly before and after the interpolation threshold.
@@
Recall that the noisy data points will be predicted based on their randomly assigned labels during this stage as the model fully interpolates the training set, we interpret the ascending $P$ curve in the over-parameterized region as a valid ‘isolation’ of the noisy labeled data in the learned feature space. When the neural network positions the learned representation of noisy data points closer to the class associated with their ‘incorrect’ labels, it results in lower $P$ prediction accuracy. Under these circumstances, similar test samples also tend to be predicted incorrectly due to the influence of these noisy labels. If the neural network successfully identifies the underlying data patterns and interpolates the correct samples around noisy data points, the prediction accuracy $P$ will be high by nearby class members. Consequently, test samples resembling the noisy training data have a higher likelihood of being accurately identified due to their proximity to the interpolated surrounding clean data points and becoming less susceptible to the influence of the ‘isolated’ noisy labeled data. This explanation accounts for the improved generalization performance for over-parameterized models.
@@ 4.2 CIFAR-10
Figure 2: The phenomenon of double descent on five-layer CNNs trained on CIFAR-10 ($N=50000$), under varying label noise ratios of $p=[0\%,10\%,20\%]$ and the prediction accuracy of noisy labeled data denoted as $P$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve of $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation generalization performance beyond context.
@@
After a baseline case of deep learning on MNIST, we extended the experiment result to image classification on CIFAR-10, using two famous neural network architectures: CNN and ResNet. We first visit the experiment when the complete CIFAR-10 dataset ($N=50000$) with zero explicit label noise ($p=0\%$) is trained with five-layer CNNs: we observe that in Figure.2(a), a proposed double descent peak arises around the interpolation threshold. We posit that the variance in observations could stem from a higher ratio of label errors in CIFAR-10 compared to MNIST, as suggested by research estimates (Zhang, 2017; Northcutt et al., 2021b). Nevertheless, when additional label noise ($p=10\%/20\%$) is introduced, the test error peak still increases accordingly, proving the positive correlation between noise and double descent according to Figure.2(b,c). While label noise increases the difficulties of models fitting the train set and shifts the interpolation threshold rightwards, the test error peak also shifts correspondingly.
@@
When we review the noisy labeled data prediction test in the learned feature space, we can still observe a correlation between the trend of generalization performance and prediction accuracy $P$. When test accuracy first increases and then decreases as the model expands, $P$ rises to about 60% and then decreases before the threshold when the peak of test error arises. Then, as the test error reduces and test accuracy rises, the $P$ curve rises again to about 60% for over-parameterized models. The observed correlation between the trend of test loss and $P$ also remains consistent. The same conclusion can be derived from the experiment context Figure.3 of ResNet18s trained on the sample task.
In conclusion, by constructing $P$ to represent the distribution of noisy labeled training data in the learned feature space, we propose that this k-NN prediction ability of the model is connected to the size of the model in the same way with double descent. We suggest that this correlation can be attributed to the model’s interpolation of correct samples around the noise labels, effectively isolating the misleading effect of the latter.
@@
Figure 3: The phenomenon of double descent on ResNet18s trained on CIFAR-10 $(N=50000)$, under varying label noise of $p=[0\%,10\%,20\%]$ and the prediction accuracy $P$ when $p>0$. Optimized with SGD for 200 epochs and decreasing learning rate. The test error curve under $p=[0\%,10\%,20\%]$ performs the double descent phenomenon and the prediction accuracy $P$ of in-context noisy data learning showed a correlation to the generalization performance beyond context.
@@ 5 Conclusion
Figure 4: Exemplification of the emergence of the double descent effect in the perspective of the present paper. Starting from the bottom of the left panel: 1. the perfect learner, in the absence of noise, will learn the perfect model containing several parameters approximately equal to the number of significant examples; 2. in the presence of noise, the perfect learner needs to learn also to distinguish information from noise and it will learn a perfect model with several parameters which is higher the number of noisy examples; 3. the imperfect learner with optimal regularization will learn a well-performing model which distinguishes information from noise using several parameters which is higher the number of examples but should not show double descent peak. On the right panel: 4. the imperfect learner with sub-optimal regularization will learn a well-performing model first interpolating noise and signal and then learning to distinguish information from noise using several parameters which is higher the number of examples showing the double descent peak; 5. the larger is the noise, the higher the double descent peak is, and the larger is the number of parameters needed to distinguish information from noise.
In this paper, we have studied the feature space of learned representations projected by various neural networks. By analyzing the behavior of the model when achieving a perfect fit on the training set, we observe that: as the model’s complexity and capacity increase, there is a clear and distinctive isolation of noisy data points within their original ‘class-mates’ in the learned feature space. This intriguing observation indicates that broader neural networks interpolate correct samples throughout the vicinity of noisy samples in the learned feature space, effectively ’isolating’ the impact of introduced label noise without explicit regularization techniques. We believe this implicit property plays a significant role in understanding how over-parameterization enhances generalization. We posit that the mechanism behind this phenomenon can be attributed to two key factors. First, the gradient descent optimization algorithm endeavors to strike a perfect balance among training data points to minimize the loss function, which can lead to the isolation of noisy data points. Second, the characteristics of high-dimensional space may facilitate a beneficial separation of noisy data points from clean data in the learned feature space.
From the perspective of this paper, the observed double decent phenomenon is associated with imperfect models, first behaving as predicted by the classical learning theory and increasing their error on the test set until they reach full interpolation on the train set. Then, they make use of the extra parameters above the interpolation point to better separate noise from information. We intentionally referred to ‘imperfect models’ because we claim that double descent should not be observable in perfect learners where perfect regularization should eliminate the effect. From our perspective, the double descent phenomenon is strictly related to imperfect models learning from noisy data.
Let’s expand on this by restricting our considerations to supervised learning for classification problems: For an ideal training set with full information (no noise and no missing cases), a learner will learn the exact model at the interpolation point where the train set is exactly classified and the error becomes zero. In this case, at interpolation, also the test set will be exactly classified because the learner has learned the exact map. Before that interpolation, the error decreases with every new example. Clearly, this system does not show any double descent. At the interpolation point, the number of model parameters is approximately equal to the number of examples in the complete train set (i.e. the model is the matching table).
In the presence of noise, the learner cannot learn the full map by interpolation. Indeed, the bias-variance trade-off predicts a U-shaped performance curve with first the learner increasing out-of-sample performances from digesting examples and then losing performances by trying to fit the in-sample noisy data. However, a learner who makes good use of regularization can do better and learn how to account for uncertainty. Likely, the number of model parameters needed to learn both the underlying map and to handle the noise is larger than the number of parameters needed to learn only the map in the ideal case without noise (the interpolation point). If the model is a perfect learner with perfect regularization, we do not expect the manifestation of any double descent, but rather a gradual improvement in out-of-sample performances with the number of parameters until a saturation point is reached where performances cannot be improved without further examples. We expect that this saturation point should occur at a number of parameters larger than the interpolation point. Double descent is instead observed for imperfect learners that at first try to interpolate the noise and then start an intrinsic regularization via over-parameterization. Our particular model of the emergence of the double descent phenomenon is illustrated In Fig. 4.
@@
Reproducibility Statement   To facilitate reproducibility, we provide a detailed description of the neural architectures in Section 3 and full training details including exact settings of hyper-parameters in Appendix A. We plan to release the source codes of our experiments to the research community.
@@ Acknowledgments
The authors acknowledge the discussion with Siddharth Chaudhary, Joe Down, Nathan D’Souza, and Huijie Yan, all UCL students who took part in the academic year 2022-23 to project COMP0031. With them, we had invaluable discussions and great inspiration. This paper originated from that project and would not have been conceived without it. Please see Chaudhary et al. (2023) for the reference to that work. The author, T.A., acknowledges partial financial support from ESRC (ES/K002309/1), EPSRC (EP/P031730/1), and EC (H2020-ICT-2018-2 825215). X.Z. knowledges partial support from the National Natural Science Foundation of China (No. 62076068).
@@ A.1 Scaling of Model Parameterization
The scaling of model parameterization to the universal layer width unit $K$ is shown in Figure.5.
@@
Figure 5: Scaling of the number of parameters as model size with layer width unit k of the three neural architectures used in our experiments. We apply a logarithmic scale to the parameter counts $P$ of all neural architectures.